OpenStudy (anonymous):

Need help checking my answer - What is the range of the following function f(x)=x^2-5 if the domain is -2, -1, 0, 1, 2? - My answer in the comments.

OpenStudy (anonymous):

Range = -5, -4, -5, -4, -5

OpenStudy (anonymous):

@Directrix Just wondering if you don't mind helping me out.

OpenStudy (anonymous):

plug each of the values of the domain in

OpenStudy (anonymous):

and then the range is the smallest and the biggest values

OpenStudy (anonymous):

I did and it turned out as the range I presented, just wondering if I got my answers right.

OpenStudy (anonymous):

let me see.

OpenStudy (anonymous):

Range = -5, -4, -5, -4, -5

OpenStudy (anonymous):

well, it seems odd the wy you got those numbers, for instance, lets plug in -2 into f(x), (-2)^2-5 4-5=-1

OpenStudy (anonymous):

you see what i mean?

OpenStudy (anonymous):

and your range should be an interval

OpenStudy (anonymous):

Ah, damn, I miscalculated. Give me a mine to re do it

OpenStudy (anonymous):

no prob : )

OpenStudy (anonymous):

did you get it?

OpenStudy (anonymous):

Range: -1, -4, -5, -4, -1

OpenStudy (anonymous):

that sounds right : ) ok so your smallest value is -5 and your biggest is -1

OpenStudy (anonymous):

so in an interval, that would be [-5,-1]

OpenStudy (anonymous):

and thats your range

OpenStudy (anonymous):

Is it possible to present a range in a set?

OpenStudy (anonymous):

Range {-1, -4, -5, -4, -1}

OpenStudy (anonymous):

mmm im not sure, if theyre telling you to do it you can

OpenStudy (anonymous):

just to be sure write it as an interval also

OpenStudy (anonymous):

Well, thanks :D

OpenStudy (anonymous):

yw :)